OpenStudy (anonymous):

which is an equation of the circle whose center is (5 -2) and whose radius is 7

OpenStudy (anonymous):

(5,-2)

OpenStudy (anonymous):

http://www.purplemath.com/modules/sqrcircle.htm

OpenStudy (anonymous):

the expression \[\sqrt{200}\] is equivalent to

OpenStudy (anonymous):

200/25 = 8 8 * 25 = 200 so sqrt(200) = sqrt(8 * 25) = sqrt( 8 *5^2) = 5*sqrt(8) = 5 * sqrt(4*2) 5*sqrt(4*2) = 5*sqrt(2^2*2) = 2*5*sqrt(2) = 10*sqrt(2)

OpenStudy (anonymous):

thank you tomo

OpenStudy (anonymous):

also sqrt(200) = sqrt(2*100) = sqrt(2*10^2) = 10*sqrt(2)

OpenStudy (anonymous):

yes it is right

OpenStudy (anonymous):

in scientific notation 54 000000 is expressed as